# Groupoid Graded Semisimple Rings

### v1

@@ 2.4. Groupoid graded modules
    Let $M$, $N$ and $P$ be $\Gamma$-graded right $R$-modules and consider homomorphisms $g\in\operatorname{HOM}_{R}(M,N)_{\gamma}$, $h\in\operatorname{HOM}_{R}(N,P)_{\sigma}$. If $\sigma\gamma$ is defined, then $h\circ g:M\to P$ is a homomorphism of degree $\sigma\gamma$, and $h\circ g=0$ otherwise.
@@
    Let $M$, $N$ and $P$ be $\Gamma$-graded left $R$-modules and consider homomorphisms $g\in\operatorname{HOM}_{R}(M,N)_{\gamma}$, $h\in\operatorname{HOM}_{R}(N,P)_{\sigma}$. If $\gamma\sigma$ is defined, then $h\circ g:M\to P$ is a homomorphism of degree $\gamma\sigma$, and $h\circ g=0$ otherwise. ∎
@@
It suffices to note that, for each $e\in\Gamma_{0}$, we have
$M(e)\neq 0\iff\mathds{1}_{e}\neq 0\iff\operatorname{END}_{R}(M)_{e}\neq 0\,.\qed$
Note that if $M$, $N$ are $\Gamma$-graded $R$-modules such that $\varphi:M\to N$ is a gr-isomorphism of modules, then $\operatorname{END}_{R}(M)\cong_{gr}\operatorname{END}_{R}(N)$ as graded rings via $g\mapsto\varphi\circ g\circ\varphi^{-1}$ for all $g\in\operatorname{END}_{R}(M)$.
@@
    via the homomorphism of additive groups $h\mapsto(h|_{M(e)})_{e\in\Gamma_{0}}$. ∎
The following result is a graded version of the generation Lemma [22, (1.1)]. If $R$ is a $\Gamma$-graded ring, $M$ is a $\Gamma$-graded $R$-module and $\{m_{i}:i\in I\}\subseteq\operatorname{h}(M)$ is a set of homogeneous generators of $M$, we will say that $\{m_{i}:i\in I\}$ is a minimal set of generators if, for each $i_{0}\in I$, $m_{i_{0}}$ does not belong to the graded submodule of $M$ generated by $\{m_{i}:i\in I\setminus\{i_{0}\}\}$.
@@
Suppose that $\{x_{j}:j\in J\}$ is another set of homogeneous generators of $M$. Each $x_{j}$ is a linear combination of $m_{i}$’s. Then, let $I_{0}$ be a minimal subset of $I$ such that $\{x_{j}:j\in J\}\subseteq\{m_{i}:i\in I_{0}\}R$. Thus, $M=\{x_{j}:j\in J\}R\subseteq\{m_{i}:i\in I_{0}\}R$ and it follows that $I_{0}=I$. If $J$ is a finite set, then $I_{0}$ also is a finite set, contradicting that $|I|$ is infinite. Therefore, $|J|$ is infinite and it follows that $|I|=|I_{0}|\leq|J|\cdot\aleph_{0}=|J|$. ∎
@@ 2.5. Conventions
In expressions like “let $M$ be a $\Gamma$-graded $R$-module” we are assuming that $M$ is a $\Gamma$-graded right $R$-module.
We will consider left modules. In this event we will make explicit the word left. If $M,N,P$ are left modules and $g\colon M\rightarrow N$, $h\colon N\rightarrow P$ are homomorphisms of modules, $g$ will act on the right. Thus, the image of $x\in M$ will be denoted by $(x)g$, and the composition $g\circ h$ means that $g$ acts first.
@@ 3.1. Pseudo-free modules
Let $M$ be a $\Gamma$-graded $R$-module. Consider a sequence of homogeneous elements $(x_{i})_{i\in I}\in\prod\limits_{i\in I}M_{\gamma_{i}}$ where $(\gamma_{i})_{i\in I}$ is a sequence of elements of $\Gamma$. A homogeneous element $x\in M_{\sigma}$ is said to be a genuine linear combination of $(x_{i})_{i\in I}$ if there exists $(a_{i})\in\bigoplus\limits_{i\in I}R_{\gamma_{i}^{-1}\sigma}$ such that $x=\sum\limits_{i\in I}x_{i}a_{i}$. We say that $(x_{i})_{i\in I}$ is pseudo-linearly independent if the only sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}1_{d(\gamma_{i})}R$ such that $\sum\limits_{i\in I}x_{i}a_{i}=0$ is $(a_{i})_{i\in I}=0$. Equivalently, if the only sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}\operatorname{h}(1_{d(\gamma_{i})}R)$ such that $\sum\limits_{i\in I}x_{i}a_{i}=0$ is $(a_{i})_{i\in I}=0$. Analogously, the only genuine linear combination of $(x_{i})_{i\in I}$ that equals $0\in M$ is with $a_{i}=0$ for all $i\in I$. The sequence $(x_{i})_{i\in I}$ is a pseudo-basis of $M$ if $(x_{i})_{i\in I}$ generates $M$ as an $R$-module and it is pseudo-linearly independent. If $M$ has a pseudo-basis, we say that $M$ is a pseudo-free module.
@@
    For each $x\in M$, there exists a unique sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}1_{d(\gamma_{i})}R$ such that
    $x=\sum_{i\in I}x_{i}a_{i}.$
@@
    Every homogenous element $x\in M$ can be uniquely expressed as a genuine linear combination of $(x_{i})_{i\in I}$. That is, for each $\sigma\in\Gamma$ and $x\in M_{\sigma}$, there exists a unique sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}R_{{\gamma_{i}}^{-1}\sigma}$ such that
    $x=\sum_{i\in I}x_{i}a_{i}.$
@@
    For any $R$-module $N$ and sequence $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N1_{d(\gamma_{i})}$, there exists a unique homomorphism of $R$-modules $f\colon M\rightarrow N$ such that $f(x_{i})=y_{i}$ for all $i\in I$.
@@
    For any $\Gamma$-graded $R$-module $N$ and sequence $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N_{\gamma_{i}}$, there exists a unique $f\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ such that $f(x_{i})=y_{i}$ for all $i\in I$.
@@
    For any $\sigma\in\Gamma$, $\Gamma$-graded $R$-module $N$ and sequence $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N_{\sigma\gamma_{i}}$, there exists a unique $f\in\operatorname{HOM}(M,N)_{\sigma}$ such that $f(x_{i})=y_{i}$ for all $i\in I$.
@@
$(1)\implies(2)$ Since $(x_{i})_{i\in I}$ generates $M$, $x=\sum\limits_{i\in I}x_{i}d_{i}$ for some $(d_{i})_{i\in I}\in\bigoplus\limits_{i\in I}R$. If we (uniquely) express each $d_{i}$ as $d_{i}=a_{i}+a^{\prime}_{i}$ where $a_{i}\in 1_{d(\gamma_{i})}R$ and $a_{i}^{\prime}\in\bigoplus\limits_{e\in\Gamma_{0}\setminus{d(\gamma_{i})}}1_{e}R$, then
@@
because $x_{i}a_{i}^{\prime}=0$ for all $i\in I$.
@@
Notice that $f(M(e))\subseteq N(e)$ and $f_{e}(M(e^{\prime}))=\{0\}$ for all $e^{\prime}\in\Delta_{0}$, $e^{\prime}\neq e$. Define also $f_{e}\colon M\rightarrow N$ as zero for all $e\in\Gamma_{0}\setminus\Delta_{0}$. Now set $f\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ as the unique $R$-module homomorphism such that $f=f_{e}$ when restricted to $M(e)$ for all $e\in\Gamma_{0}$. By Remark 2.12, (5) follows. ∎
@@
Let $M$ be a $\Gamma$-graded $R$-module. Then $M$ is a pseudo-free module if, and only if $M(e)$ is a pseudo-free module for every $e\in\Gamma_{0}$.
@@
In order to state the next result, we will say that a $\Gamma$-graded $R$-module $P$ is gr-projective if, for each $\Gamma$-graded $R$-modules $M,N$, a surjective $g\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ and $h\in\operatorname{Hom_{gr-\mathit{R}}}(P,N)$, there exists $f\in\operatorname{Hom_{gr-\mathit{R}}}(P,M)$ such that $h=gf$ [8, Proposition 35]. Now we give a direct proof of [8, Lemma 37] for object unital graded rings.
@@ 3.2. Technical results on homomorphisms of modules with degree
is well defined. Let now be $0\neq g\in\operatorname{END}_{R}(R(e))_{\gamma}$. As before, $g=m_{g(1_{e})}$, where $g(1_{e})=g(1_{e})1_{e}\in R(e)1_{e}=1_{e}R1_{e}$. ∎
@@
Following [26, Section 2.2], we define $M(\Sigma):=\bigoplus_{\sigma\in\Sigma}M(\sigma)$. When $M=R$ we write $R(\Sigma):=R_{R}(\Sigma)$. Notice that if $r(\sigma)=f$ for some $\sigma\in\Sigma$ and $f\in\Gamma_{0}\setminus\Gamma_{0}^{\prime}(M)$, then $M(\sigma)=\{0\}$. Thus, if $\Sigma^{\prime}=\{\sigma\in\Sigma\colon r(\sigma)\in\Gamma_{0}^{\prime}(M)\}$, then $M(\Sigma)=M(\Sigma^{\prime})$.
We say that $\Sigma$ is $r$-unique for $M$ if $\{r(\sigma):\sigma\in\Sigma\}\subseteq\Gamma^{\prime}_{0}(M)$ and, for each $e\in\Gamma_{0}^{\prime}(M)$, there exists at most one $\sigma\in\Sigma$ with $r(\sigma)=e$. This condition implies that, for $\sigma,\tau,\rho,\lambda\in\Sigma$, the equality $\sigma\gamma\tau^{-1}=\rho\delta\lambda^{-1}$ only holds in $\Gamma$ when
@@
The following technical result is concerned with shifts of modules and their gr-homomorphisms. It will be very useful in the sequel and some aspects of it can be regarded as an improvement of [8, Proposition 13]. The equality of Proposition 3.5(1) was given in [8, Proposition 13(a)]. We include a proof for completion.
@@
Let $M,N$ be $\Gamma$-graded $R$-modules. The following assertions hold true.
@@
$\stackrel{{\scriptstyle(*)}}{{=}}$ $\operatorname{Hom_{gr-\mathit{R}}}\left(M(\Sigma),N(\delta\gamma)\right)$
@@
$\stackrel{{\scriptstyle(*)}}{{=}}$ $\operatorname{Hom_{gr-\mathit{R}}}\left(M(\sigma\gamma^{-1}\delta^{-1}),N\right)$
@@
Notice that, if $\sigma$ or $\delta$ as before do not exist, then ()(*)( ∗ ) equals zero. Hence,

### v2

@@ 2.4. Groupoid graded modules
    Let $M$, NN and $P$ be $\Gamma$-graded right $R$-modules and consider homomorphisms $g\in\operatorname{HOM}_{R}(M,N)_{\gamma}$, $h\in\operatorname{HOM}_{R}(N,P)_{\sigma}$. If $\sigma\gamma$ is defined, then $h\circ g:M\to P$ is a homomorphism of degree $\sigma\gamma$, and $h\circ g=0$ otherwise.
@@
    Let $M$, NN and $P$ be $\Gamma$-graded left $R$-modules and consider homomorphisms $g\in\operatorname{HOM}_{R}(M,N)_{\gamma}$, $h\in\operatorname{HOM}_{R}(N,P)_{\sigma}$. If $\gamma\sigma$ is defined, then $h\circ g:M\to P$ is a homomorphism of degree $\gamma\sigma$, and $h\circ g=0$ otherwise. ∎
@@
By definition of $\Gamma^{\prime}_{0}(M)$, $e\in\Gamma^{\prime}_{0}(M)$ if and only if $M(e)\neq 0$. By definition of $\mathds{1}_{e}$, $M(e)\neq 0$ if and only if $\mathds{1}_{e}\neq 0$. By Proposition 2.10 and Lemma 2.7, $\mathds{1}_{e}\neq 0$ if and only if $\operatorname{END}_{R}(M)_{e}\neq 0$. Again by Lemma 2.7, $\operatorname{END}_{R}(M)_{e}\neq 0$ if and only if $e\in\Gamma^{\prime}_{0}(\operatorname{END}_{R}(M))$, as desired. ∎
Note that if $M$, NN are $\Gamma$-graded $R$-modules such that $\varphi:M\to N$ is a gr-isomorphism of modules, then $\operatorname{END}_{R}(M)\cong_{gr}\operatorname{END}_{R}(N)$ as graded rings via $g\mapsto\varphi\circ g\circ\varphi^{-1}$ for all $g\in\operatorname{END}_{R}(M)$.
@@
    via the homomorphism of additive groups $h\mapsto(h|_{M(e)})_{e\in\Gamma_{0}},$ where $h|_{M(e)}$ denotes the restriction of the gr-homomorphism $h\colon M\rightarrow N$ to the graded submodule $M(e)$ for each $e\in\Gamma_{0}$. Notice that $h(M(e))\subseteq N(e)$ because $h$ is a gr-homomorphism. Thus, $h|_{M(e)}\in\operatorname{Hom_{gr-\mathit{R}}}(M(e),N(e))$ for each $e\in\Gamma_{0}$. ∎
The following result is a graded version of the generation Lemma [25, (1.1)]. If $R$ is a $\Gamma$-graded ring, $M$ is a $\Gamma$-graded $R$-module and $\{m_{i}:i\in I\}\subseteq\operatorname{h}(M)$ is a set of homogeneous generators of $M$, we will say that $\{m_{i}:i\in I\}$ is a minimal set of generators if, for each $i_{0}\in I$, $m_{i_{0}}$ does not belong to the graded submodule of $M$ generated by $\{m_{i}:i\in I\setminus\{i_{0}\}\}$.
@@
Suppose that $\{x_{j}:j\in J\}$ is another set of homogeneous generators of $M$. Each $x_{j}$ is a linear combination of $m_{i}$’s. Then, let $I_{0}$ be a minimal subset of $I$ such that $\{x_{j}:j\in J\}\subseteq\{m_{i}:i\in I_{0}\}R$. Thus, $M=\{x_{j}:j\in J\}R\subseteq\{m_{i}:i\in I_{0}\}R$ and it follows that $I_{0}=I$. If $J$ is a finite set, then $I_{0}$ is also a finite set, contradicting that $|I|$ is infinite. Therefore, $|J|$ is infinite and it follows that $|I|=|I_{0}|\leq|J|\cdot\aleph_{0}=|J|$. ∎
@@ 2.5. Conventions
In expressions such as “let $M$ be a $\Gamma$-graded $R$-module,” we assume that $M$ is a $\Gamma$-graded right $R$-module.
We will consider left modules. In this event, we will make explicit the word left. If $M,N,P$ are left modules and $g\colon M\rightarrow N$, $h\colon N\rightarrow P$ are homomorphisms of modules, $g$ will act on the right. Thus, the image of $x\in M$ will be denoted by $(x)g$, and the composition $g\circ h$ means that $g$ acts first.
Let $X=\bigoplus_{\gamma\in\Gamma}X_{\gamma}$ be a $\Gamma$-graded additive group; for instance, $X$ could be a $\Gamma$-graded ring or a $\Gamma$-graded (right or left) module. When $\sigma,\tau\in\Gamma$ are such that $d(\sigma)\neq r(\tau)$ (that is, $\sigma\tau$ is not defined) we adopt the convention $X_{\sigma\tau}:=\{0\}$.
@@ 3.1. Pseudo-free modules
Let $M$ be a $\Gamma$-graded $R$-module. Consider a sequence of homogeneous elements $(x_{i})_{i\in I}\in\prod\limits_{i\in I}M_{\gamma_{i}}$ where $(\gamma_{i})_{i\in I}$ is a sequence of elements of $\Gamma$. A homogeneous element $x\in M_{\sigma}$ is said to be a genuine linear combination of $(x_{i})_{i\in I}$ if there exists $(a_{i})\in\bigoplus\limits_{i\in I}R_{\gamma_{i}^{-1}\sigma}$ such that $x=\sum\limits_{i\in I}x_{i}a_{i}$. We say that $(x_{i})_{i\in I}$ is pseudo-linearly independent if the only sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}1_{d(\gamma_{i})}R$ such that $\sum\limits_{i\in I}x_{i}a_{i}=0$ is $(a_{i})_{i\in I}=0$. Equivalently, the only sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}1_{d(\gamma_{i})}R$, with $a_{i}\in\operatorname{h}(1_{d(\gamma_{i})}R)$ for all $i\in I$, for which $\sum\limits_{i\in I}x_{i}a_{i}=0$ is $(a_{i})_{i\in I}=0$. Analogously, the only genuine linear combination of $(x_{i})_{i\in I}$ that equals $0\in M$ is with $a_{i}=0$ for all $i\in I$. The sequence $(x_{i})_{i\in I}$ is a pseudo-basis of $M$ if $(x_{i})_{i\in I}$ generates $M$ as an $R$-module and it is pseudo-linearly independent. If $M$ has a pseudo-basis, we say that $M$ is a pseudo-free module.
@@
    For each $x\in M$, there exists a unique sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}1_{d(\gamma_{i})}R$ such that x=_iI x_ia_i.
@@
    Every homogenous element $x\in M$ can be uniquely expressed as a genuine linear combination of $(x_{i})_{i\in I}$. That is, for each $\sigma\in\Gamma$ and $x\in M_{\sigma}$, there exists a unique sequence $(a_{i})_{i\in I}\in\bigoplus\limits_{i\in I}R_{{\gamma_{i}}^{-1}\sigma}$ such that x=_iI x_ia_i.
@@
    For any $R$-module NN and sequence $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N1_{d(\gamma_{i})}$, there exists a unique homomorphism of $R$-modules $f\colon M\rightarrow N$ such that $f(x_{i})=y_{i}$ for all $i\in I$.
@@
    For any $\Gamma$-graded $R$-module NN and sequence $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N_{\gamma_{i}}$, there exists a unique $f\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ such that $f(x_{i})=y_{i}$ for all $i\in I$.
@@
    For any $\sigma\in\Gamma$, $\Gamma$-graded $R$-module NN and sequence $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N_{\sigma\gamma_{i}}$, there exists a unique $f\in\operatorname{HOM}(M,N)_{\sigma}$ such that $f(x_{i})=y_{i}$ for all $i\in I$.
@@
$(1)\implies(2)$ Since $(x_{i})_{i\in I}\in\prod\limits_{i\in I}M_{\gamma_{i}}$ generates $M$, $x=\sum\limits_{i\in I}x_{i}d_{i}$ for some $(d_{i})_{i\in I}\in\bigoplus\limits_{i\in I}R$. If we (uniquely) express each $d_{i}$ as $d_{i}=a_{i}+a^{\prime}_{i}$ where $a_{i}\in 1_{d(\gamma_{i})}R$ and $a_{i}^{\prime}\in\bigoplus\limits_{e\in\Gamma_{0}\setminus{d(\gamma_{i})}}1_{e}R$, then
@@
because, for all $i\in I$, $x_{i}a_{i}^{\prime}=0$ due to the fact that $M_{\gamma_{i}}\cdot\left(\bigoplus\limits_{e\in\Gamma_{0}\setminus{d(\gamma_{i})}}1_{e}R\right)=\{0\}$.
@@
Notice that $f(M(e))\subseteq N(e)$ and $f_{e}(M(e^{\prime}))=\{0\}$ for all $^{\prime}\in\Delta_{0}, ee$$^{\prime}\neq e$. Define also $f_{e}\colon M\rightarrow N$ as zero for all $e\in\Gamma_{0}\setminus\Delta_{0}$. Now set $f\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ as the unique $R$-module homomorphism such that $f=f_{e}$ when restricted to $M(e)$ for all $e\in\Gamma_{0}$. By Remark 2.12, (5) follows. ∎
@@
Let $M$ be a $\Gamma$-graded $R$-module. Then $M$ is a pseudo-free module if and only if $M(e)$ is a pseudo-free module for every $e\in\Gamma_{0}$.
@@
In order to state the next result, we will say that a $\Gamma$-graded $R$-module $P$ is gr-projective if, for each $\Gamma$-graded $R$-modules $M,N$, a surjective $g\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ and $h\in\operatorname{Hom_{gr-\mathit{R}}}(P,N)$, there exists $f\in\operatorname{Hom_{gr-\mathit{R}}}(P,M)$ such that $h=gf$ [11, Proposition 35]. Now we give a direct proof of [11, Lemma 37] for object unital graded rings.
@@ 3.2. Technical results on homomorphisms of modules with degree
is well-defined. Let now be $0\neq g\in\operatorname{END}_{R}(R(e))_{\gamma}$. As before, $g=m_{g(1_{e})}$, where $g(1_{e})=g(1_{e})1_{e}\in R(e)1_{e}=1_{e}R1_{e}$. ∎
@@
Following [29, Section 2.2], we define $M(\Sigma):=\bigoplus_{\sigma\in\Sigma}M(\sigma)$. When $M=R$ we write $R(\Sigma):=R_{R}(\Sigma)$. Notice that if $r(\sigma)=f$ for some $\sigma\in\Sigma$ and $f\in\Gamma_{0}\setminus\Gamma_{0}^{\prime}(M)$, then $M(\sigma)=\{0\}$. Thus, if $\Sigma^{\prime}=\{\sigma\in\Sigma\colon r(\sigma)\in\Gamma_{0}^{\prime}(M)\}$, then $M(\Sigma)=M(\Sigma^{\prime})$.
We say that $\Sigma$ is $r$-unique for $M$ if $\{r(\sigma):\sigma\in\Sigma\}\subseteq\Gamma^{\prime}_{0}(M)$ and, for each $e\in\Gamma_{0}^{\prime}(M)$, there exists at most one $\sigma\in\Sigma$ with $r(\sigma)=e$. This condition implies that, for $\sigma,\tau,\rho,\lambda\in\Sigma$, the equality $\sigma\gamma\tau^{-1}=\rho\delta\lambda^{-1}$ holds in $\Gamma$ if and only if
@@
Indeed, if $\sigma\gamma\tau^{-1}=\rho\delta\lambda^{-1}$, then $r(\sigma)=r(\sigma\gamma\tau^{-1})=r(\rho\delta\lambda^{-1})=r(\rho)$. Since $\Sigma$ is $r$-unique, then $\sigma=\rho$. This implies that $\gamma\tau^{-1}=\delta\lambda^{-1}$. Thus, $r(\tau)=d(\tau^{-1})=d(\gamma\tau^{-1})=d(\delta\lambda^{-1})=d(\lambda^{-1})=r(\lambda)$. Since $\Sigma$ is $r$-unique, then $\tau=\lambda$. As a consequence, we also have $\gamma=\delta$.
@@
The following technical result is concerned with shifts of modules and their gr-homomorphisms. It will be very useful in the sequel and some aspects of it can be regarded as an improvement of [11, Proposition 13]. The equality of Proposition 3.5(1) was given in [11, Proposition 13(a)]. We include a proof for completion.
@@
Let $M,N$ be $\Gamma$-graded $R$-modules. The following assertions hold:
@@
$\mathrel{{\mathop{=}\limits^{(*)}}}$ $\operatorname{Hom_{gr-\mathit{R}}}\left(M(\Sigma),N(\delta\gamma)\right)$
@@
$\mathrel{{\mathop{=}\limits^{(*)}}}$ $\operatorname{Hom_{gr-\mathit{R}}}\left(M(\sigma\gamma^{-1}\delta^{-1}),N\right)$
@@
Notice that, if $\sigma$ or $\delta$ as before do not exist, then ()(*) equals zero. Hence,